How do I find out how many months it will take me to double my money at 5%APR? What is the formula? Thank you.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

If it's compounded monthly, the monthly interest rate = 0.05/12
-------------
2 = (1 + 0.05/12)^n (n = # of months)
log(2) = n*log(1 + 0.05/12)
n = log(2)/log(1 + 0.05/12)
n =~ 166.7 months
= 167 months

Answer by stanbon(75887)

How do I find out how many months it will take me to double my money at 5%APR? What is the formula?
2Ao = Ao(1+(r/n))^(nt)
---
Ao is the amt you invest.
r = 5%
n is the number of times per year you compound the money
t is the number of years.
---
Example: Say you compound quadterly (4 times per year).
2Ao = Ao(1+(0.05/4))^(4t)
---
2 = (1.0125)^(4t)
---
Take the log of both sides to get:
4t = log(2)/log(1.0125)
4t = 55.8
t = 13.94 years = 12*13.94 = 167.39 months
==============================================
cheers,
Stan H.
============